offset-next&assignmentProblemID 81615656 Part A 21.0 g NaHCO3 in 480.0 mL solution Submit My Answers Give Up Part B 57 o g H2SO4 in 220.0 mL solution Part C
Answer #1

molarity   = weight of substance * 1000/gram molar mass * volume of solution in ml

               = 21*1000/84*480    = 0.52M

molarity   = weight of substance * 1000/gram molar mass * volume of solution in ml

                 = 57*1000/98*220    = 2.64 M